May 19, 2017 - 200kV gun upgrade meeting

From Ciswikidb
Jump to: navigation, search

200kV Gun

  • Task and Timeline Gantt chart => media:170421_CHG.pdf
  • The electrode is out of the machine shop
    • The key was filed to prevent it from being jammed in the electrode
    • With the key fixed, the electrode assembly was measured by S&A, with and w/o the key. Danny M. compared to the model: "so the height is correct and the pitch is off 0.008 inches. those numbers seem o.k. much better than the previous numbers, and we know the key fits o.k."
  • Bubba polished the electrode with 320 grit sandpaper in prep for barrel polishing with corncob.
  • The electrode has been ultrasonic cleaned in micro in preparation for vacuum firing to 900C.

325kV Gun .... WHAT ?!?!?

media:170519_325kV_simulations.pdf

Gun2 NEG Tube

Ordered (6) 4.5" rotatable, 2-1/2" bore => (2) in hand, (4) on back order
Ordered 6" to 4-1/2" reducers =>
(3) FH sent to MMF => One will be read for NEG tube by next Friday 5/26, others week after
a) check/replace coil leads w/ 240C insulator as needed, some have old shrink wrap
b) hi pot each coil
c) measure each coil using HMIN/HMAX -2.5A/+2.5A
d) Bz v. z at r=0 ALL (and r=2,4,6 mm for ONE)
e) Bz v. I at r=0,z=0 ALL
  • BPM process : Fabricate, Clean, Leak Test, Optical Align, Electrical Acceptance and/or Gabou Line, Install
Pete Francis performing optical/electrical acceptance testing...
M20 #1 (Musson tested) - > First M20 with no serial number optically looks bad and did not pass acceptance test.
M20 #2 (Musson tested) -> Second M20 with serial no. #PM20-127 looks ok...... but probably could use an Optical Alignment
M20 #3 (Greg Marble spare) -> Third M20 tested great
M20 #4 (Greg Marble spare) -> Fourth M20 test occurred during electrical interference, re-test on Monday
  • What next?
Send to shop for modification (327755) => Request June 9
M20 Old Assembly Modification media:0990.pdf
M20 New Assembly Can media:0991.pdf
M20 New Assembly 6.00 x 4.500 reducer flangemedia:0.992.pdf
Leak check
Repeat optical/electrical acceptance
NEG coat
Gabou measurement
Install

Gun3 Removal and Diagnostics Chamber Installation

Eliminating Gun Switch Software

Software removed and interlocks tested ELOG entry
Yet to do a) remove VBV3I02 from special interlock and b) swap PLC for relays

Spellman HVPS update

Radmon update

Installed at CEBAF
Checked and ready

Summer 2017 SAD Update

media:SADCIS_v5.docx